### v2

# Study of semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays based on the SU(3) flavor symmetry

Abstract
Decays $B\to DP\ell^{+}\nu_{\ell}\leavevmode\nobreak\ (\ell=e,\mu,\tau)$ with the non-resonance, the charmed vector resonances, the charmed scalar resonances and the charmed tensor resonances are explored by using the SU(3) flavor symmetry approach. Firstly, the decay amplitudes of different modes are related by the SU(3) flavor symmetry. Then, relevant experimental data are used to constrain non-perturbative coefficients in the non-resonant and various resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Finally, using the constrained non-perturbative coefficients, the branching ratios of not-yet-measured $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonant and various charmed resonant contributions are predicted. Many branching ratios are predicted for the first time. We find that $B\to D\eta^{\prime}\ell^{+}\nu_{\ell},B_{s}\to D_{s}\eta^{\prime}\ell^{+}\nu_{\ell}$ decays only receive the non-resonant contributions, $B\to D_{s}K\ell^{+}\nu_{\ell}$, $B_{s}\to DK\ell^{+}\nu_{\ell}$, $B\to D\eta\ell^{+}\nu_{\ell}$ and $B_{s}\to D_{s}\eta\ell^{+}\nu_{\ell}$ decays receive both the non-resonant and the tensor resonant contributions, $B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ decays receive the non-resonant, the scalar resonant and the tensor resonant contributions, and other $B\to D\pi\ell^{+}\nu_{\ell}$ decays receive all kinds of contributions. These results can be used to test the SU(3) flavor symmetry approach in the four-body semileptonic $B$ decays in future experiments at LHCb and Belle-II.
I Introduction
Semileptonic $B$ decays play a key role in testing the Standard Model and understanding the heavy quark dynamics. Some three-body semileptonic $B$ decays, such as $B\to D\ell^{+}\nu_{\ell}$ and $B\to D^{*}\ell^{+}\nu_{\ell}$, have been well understood. Nevertheless, other decays, like $B\to D_{0}\ell^{+}\nu_{\ell}$, $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ and $B\to DP\ell^{+}\nu_{\ell}$ decays, received less attention. Some branching ratios of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}\leavevmode\nobreak\ (\ell^{\prime}=e,\mu)$ decays have been measured, and the experimental data from the Particle Data Group (PDG) within $2\sigma$ errors are ParticleDataGroup:2022pth
$\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ $=$ $(4.4\pm 0.8)\times 10^{-3},$ (1)
$\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{\overline{D}_{0}}$ $=$ $(2.5\pm 1.0)\times 10^{-3},$ (2)
$\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{\overline{D}^{*}_{2}}$ $=$ $(1.53\pm 0.32)\times 10^{-3},$ (3)
$\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ $=$ $(4.1\pm 1.0)\times 10^{-3},$ (4)
$\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{D_{0}}$ $=$ $(3.0\pm 2.4)\times 10^{-3},$ (5)
$\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*}_{2}}$ $=$ $(1.21\pm 0.66)\times 10^{-3},$ (6)
$\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ $=$ $(3.0^{+2.8}_{-2.4})\times 10^{-4},$ (7)
where $\mathcal{B}_{T,M}$ denote the total and $M$ resonant branching ratios. Note that, for $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ given in Eq. (1) and Eq. (4), the PDG reports results only include the $D_{0}$ and $D^{*}_{2}$ resonances, but do not include $D^{*}$ resonances. Present measurements of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays give us an opportunity to test theoretical approaches of the $B\to DP\ell^{+}\nu_{\ell}$ decays and to predict many non-measured decays, which can be further tested in the near future at LHCb and Belle-II.
Theoretically, semileptonic decays are relatively simple, since the weak and strong dynamics can be separated in these decays. All the strong dynamics in the initial and final hadrons is included in the hadronic transition form factors. The calculations of the $B\to DP$ form factors are more complex than ones of the $B\to D$ form factors or the $B\to P$ form factors Feldmann:2018kqr . So the evaluations of the $B\to DP$ form factors are difficult. In the absence of reliable calculations, the symmetry analysis can provide very useful information about the decays. SU(3) flavor symmetry is one of the popular symmetry approaches. And it has been widely used to study $b$-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn , $c$-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Qiao:2024srw ; Muller:2015lua , and light hadron decays Wang:2019alu ; Wang:2021uzi ; Xu:2020jfr ; Chang:2014iba ; Zenczykowski:2005cs ; Zenczykowski:2006se ; Cabibbo:1963yz . SU(3) flavor symmetry breaking effects due to the mass differences between the $u,d$ and $s$ quarks have also been studied, for instance, in Refs. Martinelli:2022xir ; Wang:2022fbk ; Imbeault:2011jz ; Wu:2005hi ; Dery:2020lbc ; Sasaki:2008ha ; Pham:2012db ; Geng:2018bow ; Flores-Mendieta:1998tfv ; Cheng:2012xb ; Xu:2013dta ; He:2014xha ; Yang:2015era .
Some four-body semileptonic decays $B/D\to PP\ell^{+}\nu_{\ell}$ and $B\to D^{(*)}P\ell^{+}\nu_{\ell}$ have been investigated Cheng:1993ah ; Tsai:2021ota ; Feldmann:2018kqr ; Kim:1999gm ; Kim:1998nn ; LeYaouanc:2018zyo ; Boer:2016iez ; Hambrock:2015aor ; Faller:2013dwa ; Shi:2021bvy ; Achasov:2020qfx ; Wiss:2007mr ; Wang:2016wpc ; Achasov:2021dvt ; Gustafson:2023lrz . In this work, we will explore the $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonant, the charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions by the SU(3) flavor symmetry based on the experimental data. Firstly, the hadronic amplitude relations or the form factor relations between different decay modes will be constructed. Then, the hadronic amplitudes or the form factors will be extracted by using the available data. Finally, the not-yet-measured modes will be predicted for further tests in experiments.
This paper is organized as follows. The non-resonant contributions of the $B\to DP\ell\nu_{\ell}$ decays are discussed in Sec. II. The charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions of the $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are presented in Sec. III. Finally, summary is given in Sec. IV.
II Non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
II.1 Theoretical framework
The non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are generated by $\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell}$ transition, and the effective Hamiltonian is
$\mathcal{H}_{eff}(\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell})=\frac{G_{F}}{\sqrt{2}}V_{cb}\bar{c}\gamma^{\mu}(1-\gamma_{5})b\leavevmode\nobreak\ \bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell,$ (8)
where $G_{F}$ is the Fermi constant, and $V_{cb}$ is the CKM matrix element. Decay amplitudes of the non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays can be written as
[…]
where $L_{\mu}=\bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell$ is the leptonic charged current, and $H^{\mu}=\langle D(k_{1})P(k_{2})|\bar{c}\gamma^{\mu}(1-\gamma_{5})b|B(p_{B})\rangle$ is the hadronic amplitude. Usually, $H^{\mu}$ can be obtained in terms of the form factors $F_{\bot,t,0,\|}$ of the $B\rightarrow DP$ transitions, which are are non-perturbative objects and are similar to ones of $B\to PP$ transitions Boer:2016iez . Nevertheless, the calculations of the $F_{0},F_{t},F_{\perp},F_{\parallel}$ are very difficult. In this work, the hadronic amplitude will be related by the SU(3) flavor symmetry.
Relevant meson multiplets are given first. Bottom pseudoscalar triplet $B_{i}$, charm pseudoscalar triplet $D_{i}$, charm scalar triplet $D_{0i}$, charm vector triplet $D^{*}_{i}$ and charm tensor triplet $D^{*}_{2i}$ under the SU(3) flavor symmetry of the $u$, $d$, $s$ quarks are
$B_{i}$ $=$ $\Big{(}B^{+}(\bar{b}u),B^{0}(\bar{b}d),B^{0}_{s}(\bar{b}s)\Big{)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ D_{i}=\Big{(}\overline{D}^{0}(\bar{c}u),\leavevmode\nobreak\ D^{-}(\bar{c}d),\leavevmode\nobreak\ D^{-}_{s}(\bar{c}s)\Big{)},$ (10)
$D_{0i}$ $=$ $\Big{(}\overline{D}^{0}_{0}(\bar{c}u),D^{-}_{0}(\bar{c}d),D^{-}_{s0}(\bar{c}s)\Big{)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ D^{*}_{i}=\Big{(}\overline{D}^{*0}(\bar{c}u),\leavevmode\nobreak\ D^{*-}(\bar{c}d),\leavevmode\nobreak\ D^{*-}_{s}(\bar{c}s)\Big{)},$ (11)
$D^{*}_{2i}$ $=$ $\Big{(}\overline{D}^{*0}_{2}(\bar{c}u),D^{*-}_{2}(\bar{c}d),D^{*-}_{s2}(\bar{c}s)\Big{)},$ (12)
where $i=1,2,3$ for $u$, $d$, $s$ quarks. Note that the structures of scalar $D_{0}$ mesons are not known well, and they might be four-quark states, $D_{0i}^{4q}=\Big{(}\overline{D}^{0}_{0}(\bar{c}u\bar{d}d),D^{-}_{0}(\bar{c}d\bar{u}u),D^{-}_{s0}(\bar{c}s(\bar{u}u+\bar{d}d)/\sqrt{2})\Big{)}$ Cheng:2003kg . Note that $D_{0i}$, $D^{*}_{i}$ and $D^{*}_{2i}$ will be used for the resonances in Sec. III, and they are given here together. Light pseudoscalar octets and singlets $P^{i}_{j}$ are
$P^{i}_{j}$ $=$ $\left(\begin{array}[]{cccc}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&K^{0}\\ K^{-}&\overline{K}^{0}&-\frac{2\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}\\ \end{array}\right)\,,$ (16)
with $j=1,2,3$ for $u$, $d$, $s$ quarks. The $\eta$ and $\eta^{\prime}$ are mixtures of $\eta_{1}=\frac{u\bar{u}+d\bar{d}+s\bar{s}}{\sqrt{3}}$ and $\eta_{8}=\frac{u\bar{u}+d\bar{d}-2s\bar{s}}{\sqrt{6}}$ with the mixing angle $\theta_{P}$
$\left(\begin{array}[]{c}\eta\\ \eta^{\prime}\end{array}\right)\,=\left(\begin{array}[]{cc}\mbox{cos}\theta_{P}&-\mbox{sin}\theta_{P}\\ \mbox{sin}\theta_{P}&\mbox{cos}\theta_{P}\end{array}\right)\,\left(\begin{array}[]{c}\eta_{8}\\ \eta_{1}\end{array}\right)\,.$ (23)
And $\theta_{P}=[-20^{\circ},-10^{\circ}]$ from the PDG ParticleDataGroup:2022pth will be used in our numerical analysis.
The leptonic charged current is invariant under the SU(3) flavor symmetry, and the hadronic amplitude of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decay can be parameterized by the SU(3) flavor symmetry as
$H(B\to DP)_{N}=c_{01}B^{i}P_{i}^{j}{D}_{j}+c_{02}B^{i}D_{i}P^{k}_{k},$ (24)
where $c_{01,02}$ are the non-perturbative coefficients under the SU(3) flavor symmetry. The $c_{02}$ term is suppressed by the Okubo-Zweig-Iizuka (OZI) rule Okubo:1963fa ; Lipkin:1986bi ; Lipkin:1996ny , and it only appears in the decays with $\eta,\eta^{\prime}$ final states. The idiographic hadronic amplitudes of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays are given in Tab. 1. From Tab. 1, one can see that, if ignoring the OZI suppressed $c_{02}$ contributions, all hadronic amplitudes may be related by the nonperturbative coefficient $c_{01}$.
Table 1: The hadronic amplitudes for the non-resonant $B\to DP\ell^{+}\nu$ decays under the SU(3) flavor symmetry.
   Decay modes Hadronic amplitudes    Decay modes Hadronic amplitudes
$B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ $c_{01}$ $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$ $-\frac{1}{\sqrt{2}}c_{01}$
$B^{+}\to D^{-}_{s}K^{+}\ell^{+}\nu_{\ell}$ $c_{01}$ $B^{0}\to D^{-}\eta\ell^{+}\nu_{\ell}$ $\frac{c_{01}cos\theta_{P}}{\sqrt{6}}-\frac{(c_{01}+3c_{02})sin\theta_{P}}{\sqrt{3}}$
$B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$ $\frac{1}{\sqrt{2}}c_{01}$ $B^{0}\to D^{-}\eta^{\prime}\ell^{+}\nu_{\ell}$ $\frac{c_{01}sin\theta_{P}}{\sqrt{6}}+\frac{(c_{01}+3c_{02})cos\theta_{P}}{\sqrt{3}}$
$B^{+}\to\overline{D}^{0}\eta\ell^{+}\nu_{\ell}$ $\frac{c_{01}cos\theta_{P}}{\sqrt{6}}-\frac{(c_{01}+3c_{02})sin\theta_{P}}{\sqrt{3}}$ $B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{+}\nu_{\ell}$ $c_{01}$
$B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{+}\nu_{\ell}$ $\frac{c_{01}sin\theta_{P}}{\sqrt{6}}+\frac{(c_{01}+3c_{02})cos\theta_{P}}{\sqrt{3}}$ $B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $c_{01}$
$B^{0}\to\overline{D}^{0}\pi^{-}\ell^{+}\nu_{\ell}$ $c_{01}$ $B^{0}_{s}\to D_{s}^{-}\eta\ell^{+}\nu_{\ell}$ $-\frac{2c_{01}cos\theta_{P}}{\sqrt{6}}-\frac{(c_{01}+3c_{02})sin\theta_{P}}{\sqrt{3}}$
$B^{0}\to D^{-}_{s}K^{0}\ell^{+}\nu_{\ell}$ $c_{01}$ $B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{+}\nu_{\ell}$ $-\frac{2c_{01}sin\theta_{P}}{\sqrt{6}}+\frac{(c_{01}+3c_{02})cos\theta_{P}}{\sqrt{3}}$
The differential branching ratios of the non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are Boer:2016iez
$\frac{d\mathcal{B}(B\rightarrow DP\ell^{+}\nu_{\ell})_{N}}{dq^{2}dk^{2}}=\frac{1}{2}\tau_{B}|\mathcal{N}|^{2}\beta_{\ell}(3-\beta_{\ell})|H_{N}|^{2},$ (25)
with
${|\mathcal{N}|}^{2}$ $=$ $G^{2}_{F}{|V_{cb}|}^{2}\frac{\beta_{\ell}q^{2}\sqrt{\lambda}}{3\times 2^{10}\pi^{5}m^{3}_{B}}\,,\quad\text{with}\quad\beta_{\ell}=1-\frac{m^{2}_{\ell}}{q^{2}}\,.$
$|H_{N}|^{2}$ $=$ $|F_{0}|^{2}+\frac{2}{3}(|F_{\parallel}|^{2}+|F_{\perp}|^{2})+\frac{3m_{\ell}^{2}}{q^{2}(3-\beta_{\ell})}|F_{t}|^{2},$ (26)
where $\tau_{M}$($m_{M}$) is lifetime(mass) of $M$ particle. The ranges of integration are given by $(m_{D}+m_{P})^{2}\leq k^{2}\leq(m_{B}-m_{\ell})^{2}$ and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-\sqrt{k^{2}})^{2}$. If we ignore $|F_{t}|^{2}$ term since it is proportional to $m_{\ell}^{2}$ and it is small when $\ell=e,\mu$, $|H_{N}|^{2}$ is only include the hadronic part. Noted that although $|F_{t}|^{2}$ term might be large when $\ell=\tau$, it is difficult to estimate its contribution in this work, so we still ignore it. Then $H_{N}$, which only includes hadronic part, follow the relationship of the SU(3) flavor symmetry in Tab. 1.
II.2 Numerical results
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, no any process has been measured until now. However, as given in Eq. (7), $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ has been measured. The $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ mode can decay via the non-resonance and the $D^{*}_{2}$ tensor meson resonance. In the subsequent analysis in Sec. III, the contributions of $D^{*}_{2}$ tensor meson resonance are far less than the experimental data given in Eq. (7). So we think that the non-resonant contributions are dominant in the $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, $i.e.$, $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}\approx\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$. The experimental data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ are used to determine $c_{01}$ in the non-resonant $B\to DP\ell^{+}\nu_{\ell^{\prime}}$ decays (Due to poor relevant experimental data, the OZI suppressed $c_{02}$ term is ignored). Then many other branching ratios of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays can be predicted by using the data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$, which are listed in the second column of Tab. 2.
From the second column of Tab. 2, one can see that many branching ratio central values of the non-resonant $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, such as $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Nevertheless, other decays, for example, the non-resonant $B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}\to D^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, and all $B\to DP\tau^{+}\nu_{\tau}$ decays, are strongly suppressed by the narrow phase spaces, their branching ratio central values are on the orders of $\mathcal{O}(10^{-5}-10^{-7})$, and they might not be observed by the experiments in the near future.
Table 2: Branching ratios for the $B\to DP\ell^{+}\nu_{\ell}$ decays within $2\sigma$ errors (in units of $10^{-3}$). $\mathcal{B}_{N}$ denotes the non-resonant branching ratios, $\mathcal{B}_{[R]}$ denotes the $R$ resonant ones, edenotes experimental data within $2\sigma$ errors, and denotes the results obtained by considering the resonance width effects.
     Decay modes $\mathcal{B}_{N}$          $\mathcal{B}_{[D^{*}]}$         $\mathcal{B}_{[D_{0}]}$                $\mathcal{B}_{[D^{*}_{2}]}$
$B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.64\pm 0.52$cdots ${}^{2.64\pm 0.86_{[D^{0}_{0}]}}_{2.5\pm 1.0^{e}}$ ${}^{1.33\pm 0.12_{[D^{*0}_{2}]}}_{1.53\pm 0.32^{e}}$
$B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ ${}^{0.32\pm 0.26}_{0.30^{+0.28e}_{-0.24}}$cdots $\cdots$ $[5.66\times 10^{-15},2.15\times 10^{-7}]$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.32\pm 0.26$ ${}^{35.10\pm 2.68_{[D^{*0}]}}_{31.22\pm 2.25^{\dagger}_{[D^{*0}]}}$ $1.34\pm 0.44_{[D^{0}_{0}]}$ $0.70\pm 0.06$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.11\pm 0.10$cdots $\cdots$ $(4.36\pm 1.22)\times 10^{-3}$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.038\pm 0.033$cdots $\cdots$ $\cdots$
$B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.60\pm 0.49$ ${}^{33.67\pm 2.11_{[D^{*-}]}}_{30.14\pm 1.72^{\dagger}_{[D^{*-}]}}$ ${}^{2.48\pm 0.81_{[D^{-}_{0}]}}_{3.0\pm 2.4^{e}}$ ${}^{1.28\pm 0.13_{[D^{*-}_{2}]}}_{1.21\pm 0.66^{e}}$
$B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.30\pm 0.24$cdots $\cdots$ $[4.07\times 10^{-14},5.30\times 10^{-6}]$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.30\pm 0.24$ ${}^{15.29\pm 1.21_{[D^{*-}]}}_{13.69\pm 1.05^{\dagger}_{[D^{*-}]}}$ $1.23\pm 0.40_{[D^{-}_{0}]}$ $0.62\pm 0.06$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.10\pm 0.09$cdots $\cdots$ $(3.54\pm 1.40)\times 10^{-3}$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.035\pm 0.031$cdots $\cdots$ $\cdots$
$B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.41\pm 0.33$cdots $\cdots$ $1.23\pm 0.18$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.40\pm 0.33$cdots $\cdots$ $1.11\pm 0.16$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.15\pm 0.13$cdots $\cdots$ $(1.67\pm 0.59)\times 10^{-2}$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$ $0.095\pm 0.081$cdots $\cdots$ $\cdots$
$B^{+}\to D^{-}\pi^{+}\tau^{+}\nu_{\tau}$ $0.091\pm 0.074$cdots $0.35\pm 0.12_{[D^{0}_{0}]}$ $(8.62\pm 1.82)\times 10^{-2}$${}_{[D^{*0}_{2}]}$
$B^{+}\to D^{-}_{s}K^{+}\tau^{+}\nu_{\tau}$ $0.022\pm 0.017$cdots $\cdots$ $[3.54\times 10^{-16},1.45\times 10^{-8}]$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\pi^{0}\tau^{+}\nu_{\tau}$ $0.047\pm 0.038$ ${}^{8.59\pm 0.66_{[D^{*0}]}}_{7.63\pm 0.55^{\dagger}_{[D^{*0}]}}$ $0.18\pm 0.06_{[D^{0}_{0}]}$ $(4.52\pm 0.95)\times 10^{-2}$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\eta\tau^{+}\nu_{\tau}$ $0.0085\pm 0.0072$cdots $\cdots$ $(2.81\pm 1.00)\times 10^{-4}$${}_{[D^{*0}_{2}]}$
$B^{+}\to\overline{D}^{0}\eta^{\prime}\tau^{+}\nu_{\tau}$ $0.00086\pm 0.00074$cdots $\cdots$ $\cdots$
$B^{0}\to\overline{D}^{0}\pi^{-}\tau^{+}\nu_{\tau}$ $0.086\pm 0.070$ ${}^{8.22\pm 0.52_{[D^{*-}]}}_{7.34\pm 0.42^{\dagger}_{[D^{*-}]}}$ $0.32\pm 0.11_{[D^{-}_{0}]}$ $(8.34\pm 1.78)\times 10^{-2}$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}_{s}K^{0}\tau^{+}\nu_{\tau}$ $0.020\pm 0.016$cdots $\cdots$ $[2.51\times 10^{-15},3.20\times 10^{-7}]$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\pi^{0}\tau^{+}\nu_{\tau}$ $0.043\pm 0.035$ ${}^{3.73\pm 0.30_{[D^{*-}]}}_{3.34\pm 0.26^{\dagger}_{[D^{*-}]}}$ $0.16\pm 0.06_{[D^{-}_{0}]}$ $(4.06\pm 0.87)\times 10^{-2}$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\eta\tau^{+}\nu_{\tau}$ $0.0077\pm 0.0066$cdots $\cdots$ $(2.34\pm 1.03)\times 10^{-4}$${}_{[D^{*-}_{2}]}$
$B^{0}\to D^{-}\eta^{\prime}\tau^{+}\nu_{\tau}$ $0.00077\pm 0.00067$cdots $\cdots$ $\cdots$
$B^{0}_{s}\to\overline{D}^{0}K^{-}\tau^{+}\nu_{\tau}$ $0.040\pm 0.033$cdots $\cdots$ $(7.86\pm 1.99)\times 10^{-2}$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D^{-}\overline{K}^{0}\tau^{+}\nu_{\tau}$ $0.039\pm 0.032$cdots $\cdots$ $(7.10\pm 1.79)\times 10^{-2}$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D_{s}^{-}\eta\tau^{+}\nu_{\tau}$ $0.011\pm 0.009$cdots $\cdots$ $(1.04\pm 0.44)\times 10^{-3}$${}_{[D^{*-}_{s2}]}$
$B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\tau^{+}\nu_{\tau}$ $0.0020\pm 0.0017$cdots $\cdots$ $\cdots$
III Decays $B\to DP\ell^{+}\nu_{\ell}$ with the $D^{*},D_{0},D^{*}_{2}$ resonances
For the $B\to DP\ell^{+}\nu_{\ell}$ decays with the resonances, when the decay widths of the resonance states are very narrow, the resonant branching ratios respect a simple factorization relation
$\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell},D_{J}\rightarrow DP)=\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})\times\mathcal{B}(D_{J}\rightarrow DP),$ (27)
due to parity conservation, only $D_{J}=D_{0}/D^{*}/D^{*}_{2}$ resonances are considered for the $B\to DP\ell^{+}\nu_{\ell}$ decays Kim:1998np . And this result is also a good approximation for wider resonances. Eq. (27) will be used in our analysis for resonant $B\rightarrow D_{J}(\rightarrow DP)\ell^{+}\nu_{\ell}$ decays. We first calculate $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{J}\rightarrow DP)$ by the SU(3) flavor symmetry, then obtain $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell},D_{J}\rightarrow DP)$. The width effects of some resonance states will be analyzed later.
III.1 Decays $B\to D_{J}\ell^{+}\nu_{\ell}$
The effective Hamiltonian is given in Eq. (8), and the amplitudes of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays also can be factorized in the product of the matrix elements of leptonic and hadronic currents. The differential branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are Ivanov:2019nqd
$\frac{d\mathcal{B}(B\to{D_{J}}\ell^{+}\nu_{\ell})}{dq^{2}}=\frac{\tau_{B}G_{F}^{2}|V_{cb}|^{2}\lambda^{1/2}(q^{2}-m_{\ell}^{2})^{2}}{24(2\pi)^{3}m_{B}^{3}q^{2}}{\cal H}_{\rm total},$ (28)
with
${\cal H}_{\rm total}$ $=$ $({\cal H}_{U}+{\cal H}_{L})\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)+\frac{3m_{\ell}^{2}}{2q^{2}}{\cal H}_{S},$ (29)
${\cal H}_{U}$ $=$ $|H^{D_{J}}_{+}|^{2}+|H^{D_{J}}_{-}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{L}=|H^{D_{J}}_{0}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{S}=|H^{D_{J}}_{t}|^{2},$ (30)
where $\lambda\equiv\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$, and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-m_{D_{J}})^{2}$. The hadronic helicity amplitudes are defined through the hadronic matrix elements
$H^{D_{J}}_{m^{\prime}}$ $=$ $\epsilon^{*}_{\beta}(m^{\prime})\langle{D_{J}}(p,\varepsilon^{*})|\bar{c}\gamma^{\beta}(1-\gamma_{5})b|B(p_{B})\rangle,$ (31)
where $\varepsilon^{*}$ is the polarization of the $D^{*}$ and $D^{*}_{2}$ mesons, and $\epsilon_{\mu}(m)$ is the polarization of the virtual $W$ with $m=0,t,\pm 1$.
The hadronic helicity amplitudes related to the form factors are
$H^{D}_{\pm}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ H^{D}_{0}=\frac{2m_{B}|\vec{p}_{D}|}{\sqrt{q^{2}}}f^{BD}_{+}(q^{2}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ H^{D}_{t}=\frac{m_{B}^{2}-m_{D}^{2}}{\sqrt{q^{2}}}f^{BD}_{0}(q^{2}),$ (32)
for $B\to D\ell^{+}\nu_{\ell}$ decays,
$H^{D_{0}}_{\pm}$ $=$ $0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ H^{D_{0}}_{0}=\frac{i2m_{B}|\vec{p}_{D_{0}}|}{\sqrt{q^{2}}}f^{BD_{0}}_{1}(q^{2}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ H^{D_{0}}_{t}=\frac{i(m_{B}^{2}-m_{D_{0}}^{2})}{\sqrt{q^{2}}}f^{BD_{0}}_{0}(q^{2}),$ (33)
for $B\to D_{0}\ell^{+}\nu_{\ell}$ decays,
$H^{D^{*}}_{\pm}$ $=$ $(m_{B}+m_{D^{*}})A^{BD^{*}}_{1}(q^{2})\mp\frac{2m_{B}|\vec{p}_{D^{*}}|}{(m_{B}+m_{D^{*}})}V^{BD^{*}}(q^{2}),$ (34)
$H^{D^{*}}_{0}$ $=$ $\frac{1}{2m_{D^{*}}\sqrt{q^{2}}}\left[(m_{B}^{2}-m_{D^{*}}^{2}-q^{2})(m_{B}+m_{D^{*}})A^{BD^{*}}_{1}(q^{2})-\frac{4m_{B}^{2}|\vec{p}_{D^{*}}|^{2}}{m_{B}+m_{D^{*}}}A^{BD^{*}}_{2}(q^{2})\right],$ (35)
$H^{D^{*}}_{t}$ $=$ $\frac{2m_{B}|\vec{p}_{D^{*}}|}{\sqrt{q^{2}}}A^{BD^{*}}_{0}(q^{2}),$ (36)
for $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, and
$H^{D^{*}_{2}}_{\pm}$ $=$ $\frac{2|\vec{p}_{D^{*}_{2}}|}{\sqrt{6}m_{D^{*}_{2}}}\left[(m_{B}+m_{D^{*}_{2}})A^{BD^{*}_{2}}_{1}(q^{2})\mp\frac{2m_{B}|\vec{p}_{D^{*}_{2}}|}{(m_{B}+m_{D^{*}_{2}})}V^{BD^{*}_{2}}(q^{2})\right],$ (37)
$H^{D^{*}_{2}}_{0}$ $=$ $\frac{|\vec{p}_{D^{*}_{2}}|}{\sqrt{2}m_{D^{*}_{2}}}\frac{1}{2m_{D^{*}_{2}}\sqrt{q^{2}}}\left[(m_{B}^{2}-m_{D^{*}_{2}}^{2}-q^{2})(m_{B}+m_{D^{*}_{2}})A^{BD^{*}_{2}}_{1}(q^{2})-\frac{4m_{B}^{2}|\vec{p}_{D^{*}_{2}}|^{2}}{m_{B}+m_{D^{*}_{2}}}A^{BD^{*}_{2}}_{2}(q^{2})\right],$ (38)
$H^{D^{*}_{2}}_{t}$ $=$ $\frac{|\vec{p}_{D^{*}_{2}}|}{\sqrt{2}m_{D^{*}_{2}}}\frac{2m_{B}|\vec{p}_{D^{*}_{2}}|}{\sqrt{q^{2}}}A^{BD^{*}_{2}}_{0}(q^{2}),$ (39)
for $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, where $|\vec{p}_{D_{J}}|\equiv\sqrt{\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})}/(2m_{B})$.
The form factors of the $B\to D_{J}$ transitions are defined as Cheng:2017pcq ; Cheng:2003sm
$D(p)\left|\bar{c}\gamma_{\mu}b\right|B(p_{B})\right>$ $=$ $f_{1}^{BD}(q^{2})\Big{(}(p+p_{B})_{\mu}-\frac{m_{B}^{2}-m_{D}^{2}}{q^{2}}q_{\mu}\Big{)}+f_{0}^{BD}(q^{2})\frac{m_{B}^{2}-m_{D}^{2}}{q^{2}}q_{\mu},$ (40)
$D_{0}(p)\left|\bar{c}\gamma_{\mu}\gamma_{5}b\right|B(p_{B})\right>$ $=$ $-i\Bigg{[}f_{1}^{BD_{0}}(q^{2})\Big{(}(p+p_{B})_{\mu}-\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Big{)}+f_{0}^{BD_{0}}(q^{2})\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Bigg{]},$ (41)
$D^{*}(p,\varepsilon^{*})\left|\bar{c}\gamma_{\mu}(1-\gamma_{5})b\right|B(p_{B})\right>$ $=$ $\frac{2V^{BD^{*}}(q^{2})}{m_{B}+m_{D^{*}}}\epsilon_{\mu\nu\alpha\beta}\varepsilon^{*\nu}p^{\alpha}_{B}p^{\beta}$ (42)
$-i\left[\varepsilon^{*}_{\mu}(m_{B}+m_{D^{*}})A^{BD^{*}}_{1}(q^{2})-(p_{B}+p)_{\mu}(\varepsilon^{*}.p_{B})\frac{A^{BD^{*}}_{2}(q^{2})}{m_{B}+m_{D^{*}}}\right]$
$+iq_{\mu}(\varepsilon^{*}.p_{B})\frac{2m_{D^{*}}}{q^{2}}\Big{[}A^{BD^{*}}_{3}(q^{2})-A^{BD^{*}}_{0}(q^{2})\Big{]},$
$D^{*}_{2}(p,\varepsilon^{*})\left|\bar{c}\gamma_{\mu}(1-\gamma_{5})b\right|B(p_{B})\right>$ $=$ $\frac{2iV^{BD^{*}_{2}}(q^{2})}{m_{B}+m_{D^{*}_{2}}}\epsilon_{\mu\nu\alpha\beta}e^{*\nu}p^{\alpha}_{B}p^{\beta}$ (43)
$+2m_{D^{*}_{2}}\frac{e^{*}\cdot q}{q^{2}}q_{\mu}A^{BD^{*}_{2}}_{0}(q^{2})+(m_{B}+m_{D^{*}_{2}})\Big{(}e^{*}_{\mu}-\frac{e^{*}\cdot q}{q^{2}}q_{\mu}\Big{)}A^{BD^{*}_{2}}_{1}(q^{2})$
$-\frac{e^{*}\cdot q}{m_{B}+m_{D^{*}_{2}}}\Big{(}(p_{B}+p)_{\mu}-\frac{m_{B}^{2}-m^{2}_{D^{*}_{2}}}{q^{2}}q_{\mu}\Big{)}A^{BD^{*}_{2}}_{2}(q^{2}),$
where $s=q^{2}$ ($q=p_{B}-p$) and $e^{*\nu}\equiv\frac{\varepsilon^{*\mu\nu}\cdot p_{B\mu}}{m_{B}}$.
Now one can obtain the branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays by the relevant form factors, which depend on the different methods. In this work, we use the SU(3) flavor symmetry to obtain the relations of the hadronic amplitudes, and the same relations are also true for the form factors. In terms of the SU(3) flavor symmetry, the hadronic helicity amplitudes defined in Eq. (31) can be parameterized as
$H^{D_{J}}_{m^{\prime}}=C^{D_{J}}_{0}B_{i}(D_{J})^{i},$ (44)
where $C^{D_{J}}_{0}$ are the non-perturbative coefficients under the SU(3) flavor symmetry. For the charmed four-quark states $D_{0}^{4q}$, $H^{D_{0}}_{m^{\prime}}=C^{4q,D_{0}}_{0}B_{i}(D_{J})^{ij}_{j}$. And the hadronic amplitude relations for the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are summarized in Tab. 3.
Table 3: The hadronic amplitudes for $B\to D_{J}\ell^{+}\nu$ decays under the SU(3) flavor symmetry.
Decay moeds SU(3) hadronic amplitudes Decay moeds SU(3) hadronic amplitudes
$B^{+}\to\overline{D}^{0}\ell^{+}\nu_{\ell}$ $C^{D}_{0}$ $B^{+}\to\overline{D}^{*0}\ell^{+}\nu_{\ell}$ $C^{D^{*}}_{0}$
$B^{0}\to D^{-}\ell^{+}\nu_{\ell}$ $C^{D}_{0}$ $B^{0}\to D^{*-}\ell^{+}\nu_{\ell}$ $C^{D^{*}}_{0}$
$B^{0}_{s}\to D^{-}_{s}\ell^{+}\nu_{\ell}$ $C^{D}_{0}$ $B^{0}_{s}\to D^{*-}_{s}\ell^{+}\nu_{\ell}$ $C^{D^{*}}_{0}$
$B^{+}\to\overline{D}^{0}_{0}\ell^{+}\nu_{\ell}$ $C^{D_{0}}_{0},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C^{4q,D_{0}}_{0}$ $B^{+}\to\overline{D}^{*0}_{2}\ell^{+}\nu_{\ell}$ $C^{D^{*}_{2}}_{0}$
$B^{0}\to D^{-}_{0}\ell^{+}\nu_{\ell}$ $C^{D_{0}}_{0},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C^{4q,D_{0}}_{0}$ $B^{0}\to D^{*-}_{2}\ell^{+}\nu_{\ell}$ $C^{D^{*}_{2}}_{0}$
$B^{0}_{s}\to D^{-}_{s0}\ell^{+}\nu_{\ell}$ $C^{D_{0}}_{0},\leavevmode\nobreak\ \leavevmode\nobreak\ \sqrt{2}C^{4q,D_{0}}_{0}$ $B^{0}_{s}\to D^{*-}_{s2}\ell^{+}\nu_{\ell}$ $C^{D^{*}_{2}}_{0}$
The relations in Tab. 3 will be used for the form factors $F(0)$, which are $f^{BD}_{1}(0)$ in the $B\to D\ell^{+}\nu_{\ell}$ decays, $V^{BD^{*}}(0)$ in the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, $f^{BD_{0}}_{1}(0)$ in the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays, and $A^{BD^{*}_{2}}_{1}(0)$ in $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays. The form factors $F(0)$ are determined by the relevant experimental data. Other form factors $F_{i}(0)$ can be expressed as $r_{i}\times F(0)$, and the values of the ratios $r_{i}=\frac{F_{i}(0)}{F(0)}$ are taken from Ref. Totten:2000ab for the $B\to D/D^{*}\ell^{+}\nu_{\ell}$ decays, from Ref. Verma:2011yw for the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays and from Ref. Chen:2021ywv for the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays. Taking the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays as an example, there are four form factors $V^{BD^{*}}(0)$ and $A^{BD^{*}}_{0,1,2}(0)$ in the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, $A^{BD^{*}}_{0,1,2}(0)$ are expressed by $r_{0,1,2}\times V^{BD^{*}}(0)$, and the values of $r_{0,1,2}=\frac{A^{BD^{*}}_{0,1,2}(0)}{V^{BD^{*}}(0)}$ are taken from Ref. Totten:2000ab , and then there is only one parameter $V^{BD^{*}}(0)$ in the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, and it can be determined by the experimental data of the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays.
Now we give our branching ratio predictions of the semileptonic $B\to D_{J}\ell^{+}\nu_{\ell}$ decays under the SU(3) flavor symmetry. If not specially specified, the theoretical input parameters, such as the lifetimes, the masses, and the experimental data within the $2\sigma$ error bars from PDG ParticleDataGroup:2022pth will be used in our numerical analysis.
Theoretically, exclusive semileptonic $B\to D/D^{*}\ell^{+}\nu_{\ell}$ are well understood. Although the $B\to D\ell^{+}\nu_{\ell}$ decays are not used for the four-body semileptonic decay branching ratios, there are five experimental data in the $B\to D\ell^{+}\nu_{\ell}$ decays, which could be used to test the SU(3) flavor symmetry approach, so we present their results here. The experimental data of the $B\to D\ell^{+}\nu_{\ell}$ decays are listed in the second column of Tab. 5, which are used to constrain the only one free parameter $f^{BD}_{1}(0)$. We obtain that $f^{BD}_{1}(0)=0.66\pm 0.05$, which agrees with $0.67$ given in Ref. Totten:2000ab . Then one can predict the branching ratios of the $B^{0}_{s}\to D^{-}_{s}\ell^{+}\nu_{\ell}$ decays in terms of the constrained $f^{BD}_{1}(0)$, which are listed in the third column of Tab. 5.
Table 4: The experimental data and the SU(3) flavor symmetry predictions of $\mathcal{B}(B\to D/D^{*}\ell^{+}\nu_{\ell})$ within $2\sigma$ errors (in units of $10^{-2}$). adenotes that the experimental data are not used to constrain the parameter $C^{D^{*}}_{0}$.
Branching ratios Exp. data ParticleDataGroup:2022pth Our predictions Branching ratios Exp. data ParticleDataGroup:2022pth Our predictions
$\mathcal{B}(B^{+}\to\overline{D}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.30\pm 0.18$ $2.34\pm 0.14$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $5.58\pm 0.44$ $5.41\pm 0.27$
$\mathcal{B}(B^{0}\to D^{-}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.24\pm 0.18$ $2.19\pm 0.13$ $\mathcal{B}(B^{0}\to D^{*-}\ell^{\prime+}\nu_{\ell^{\prime}})$ $4.97\pm 0.24$ $4.97\pm 0.24$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $2.20\pm 0.14$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $4.99\pm 0.28$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\mu^{+}\nu_{\mu})$ $2.44\pm 0.46$ $2.20\pm 0.14$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\mu^{+}\nu_{\mu})$ $5.30\pm 1.0$ $4.98\pm 0.28$
$\mathcal{B}(B^{+}\to\overline{D}^{0}\tau^{+}\nu_{\tau})$ $0.77\pm 0.50$ $0.68\pm 0.04$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}\tau^{+}\nu_{\tau})$ $1.88\pm 0.40^{a}$ $1.35\pm 0.07$
$\mathcal{B}(B^{0}\to D^{-}\tau^{+}\nu_{\tau})$ $1.05\pm 0.46$ $0.64\pm 0.04$ $\mathcal{B}(B^{0}\to D^{*-}\tau^{+}\nu_{\tau})$ $1.58\pm 0.18^{a}$ $1.21\pm 0.06$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\tau^{+}\nu_{\tau})$cdots $0.63\pm 0.04$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\tau^{+}\nu_{\tau})$cdots $1.20\pm 0.07$
Table 5: The experimental data and the SU(3) flavor symmetry predictions of $\mathcal{B}(B\to D_{0}/D_{2}^{*}\ell^{+}\nu_{\ell})$ within $2\sigma$ errors. 2q(4q)denote the two(four) quark state predictions. $\mathcal{B}(B\to D_{0}/D^{*}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ is in unit of $10^{-3}$, and $\mathcal{B}(B\to D_{0}/D^{*}_{2}\tau^{+}\nu_{\tau})$ is in unit of $10^{-4}$.
Branching ratios Exp. data Belle-II:2022evt Our predictions Branching ratios Exp. data Belle-II:2022evt Our predictions
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $4.2\pm 1.6$ $3.98\pm 1.30$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.9\pm 0.6$ $3.20\pm 0.30$
$\mathcal{B}(B^{0}\to D^{-}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $3.9\pm 1.4$ $3.71\pm 1.21$ $\mathcal{B}(B^{0}\to D^{*-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.7\pm 0.6$ $2.99\pm 0.29$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $4.43\pm 1.54^{2q}$$8.84\pm 3.08^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $2.72\pm 0.27$
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\tau^{+}\nu_{\tau})$cdots $5.23\pm 1.85$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\tau^{+}\nu_{\tau})$cdots $2.15\pm 0.47$
$\mathcal{B}(B^{0}\to D^{-}_{0}\tau^{+}\nu_{\tau})$cdots $4.86\pm 1.72$ $\mathcal{B}(B^{0}\to D^{*-}_{2}\tau^{+}\nu_{\tau})$cdots $1.97\pm 0.43$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\tau^{+}\nu_{\tau})$cdots $6.75\pm 2.40^{2q}$$13.50\pm 4.80^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})$cdots $1.73\pm 0.38$
For the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, there are also five measured modes, and they are listed in the fifth column of Tab. 5. $\mathcal{B}(B^{+}\to\overline{D}^{*0}\tau^{+}\nu_{\tau})$ and $\mathcal{B}(B^{0}\to D^{*-}\tau^{+}\nu_{\tau})$ are not used to constrain the only free parameter $V^{BD^{*}}(0)$. We obtain that $V^{BD^{*}}(0)=0.65\pm 0.05$ from three experimental data of $\mathcal{B}(B_{(s)}\to D^{*}\ell^{\prime+}\nu_{\ell^{\prime}})$, which is smaller than $0.76$ given in Ref. Totten:2000ab . Then one can predict other branching ratios of the $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, which are listed in the last column of Tab. 5. One can see that our prediction and experimental data of $\mathcal{B}(B^{+}\to\overline{D}^{*0}\tau^{+}\nu_{\tau})$ intersect within $2\sigma$ error ranges, nevertheless, our prediction of $\mathcal{B}(B^{0}\to D^{*-}\tau^{+}\nu_{\tau})$ is slightly smaller than its data, and they will agree within $3\sigma$ error ranges.
For the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays, only two decay modes have been measured, and they are listed in the second column of Tab. 5, which are used to constrain the parameter $f^{BD_{0}}_{1}(0)$. Our constrained $f^{BD_{0}}_{1}(0)=0.38\pm 0.09$, which is obviously larger than $0.27\pm 0.03$ given in Ref. Verma:2011yw . Our branching ratio predictions of the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays are listed in the third column of Tab. 5. The branching ratio predictions of $B^{0}_{s}\to D^{-}_{s0}\ell^{+}\nu_{\ell}$ are different between the two quark state and the four quark state, the prediction with four quark state are 2 times of one with the two quark state. In the later analysis of $B\to DP\ell\nu_{\ell}$ with $D_{0}$ resonances, the results of the two quark state will be used.
For the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, only $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ and $\mathcal{B}(B^{0}\to D^{*-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ have been measured, and they are listed in the fifth column of Tab. 5. We obtain $A_{1}^{BD^{*}_{2}}(0)=0.46\pm 0.06$ from two measured branching ratios, which are consistent with $0.63^{+0.11}_{-0.12}$ within $2\sigma$ errors given in Ref. Chen:2021ywv . The branching ratio predictions of the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays are listed in the last column of Tab. 5. Decays $B^{0}_{s}\to D^{*-}_{s2}\ell^{+}\nu_{\ell}$ have been calculated by the QCD sum rule approach for different scale parameter $\mu=2/3/4$ GeV Azizi:2014nta , for examples, $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})=(3.07\pm 1.40)\times 10^{-3}$ and $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})=(1.03\pm 0.61)\times 10^{-3}$ for $\mu=4$ GeV. The predictions of $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$ in Ref. Azizi:2014nta are consistent with ours, nevertheless, the predictions of $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})$ in Ref. Azizi:2014nta are smaller than ours.
Until now, most of the SU(3) flavor symmetry predictions of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are quite coincident with their experimental data within $2\sigma$ errors. The SU(3) flavor breaking effects mainly come from different masses of $u$, $d$, and $s$ quarks. Since $m_{u,d}$ are much smaller than $m_{s}$, the SU(3) breaking effects due to a non-zero $m_{s}$ dominate the SU(3) breaking effects He:2014xha . If considering the SU(3) flavor breaking effects due to a non-zero $m_{s}$, the non-perturbative coefficients of the $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}/D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are different from those of the $B^{+}\to\overline{D}^{0}/\overline{D}^{*0}/\overline{D}^{0}_{0}/\overline{D}^{*0}_{2}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}/D^{*-}/D^{-}_{0}/D^{*-}_{2}\ell^{+}\nu_{\ell}$ decays. As given in Tab. 5, decays $B^{0}_{s}\to D^{-}_{s}\mu^{+}\nu_{\mu}$ and $B^{0}_{s}\to D^{*-}_{s}\mu^{+}\nu_{\mu}$ have been measured, so one can estimate the SU(3) flavor breaking effects due to a non-zero $m_{s}$ in the $B\to D/D^{*}\ell^{+}\nu_{\ell}$ decays. Comparing our SU(3) flavor symmetry predictions and their experimental measurements of $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}\mu^{+}\nu_{\mu}$ decays within $2\sigma$ errors, one can find the SU(3) breaking contributions to $\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\ell^{+}\nu_{\ell})$ should be less than roughly 23% and 20% of their experimental central values, respectively. After the $B^{0}_{s}\to D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are measured, one can estimate the SU(3) flavor breaking effects in the $B\to D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays.
III.2 Decays $D_{J}\to DP$
Non-leptonic two-body $D_{J}\to DP$ decays via strong or electromagnetic interaction are presented in this section. In terms of the SU(3) flavor symmetry, the decay amplitudes of the strong or electromagnetic $D_{J}\to DP$ decays can be parameterized as
$A(D_{J}\to DP)=a^{D_{J}}_{01}(D_{J})_{i}P^{i}_{j}D^{j}+a^{D_{J}}_{02}(D_{J})_{i}D^{i}P^{j}_{j},$ (45)
where $a^{D_{J}}_{01,02}$ are the non-perturbative coefficients, and all $D_{J}$ are two-quark states. $a^{D_{J}}_{02}$ are OZI suppressed and it will be ignored in later numerical analysis. The decay amplitudes for each $D^{*}_{2}\to DP$ decay are summarized in Tab. 7. The decay amplitudes for $D^{*}/D_{0}\to DP$ can be obtained by replacing $a^{D^{*}_{2}}_{01,02}$ listed in Tab. 7 with $a^{D^{*},D_{0}}_{01,02}$ only if their phase spaces are allowed.
Then the branching ratios of the $D_{J}\to DP$ decays can be written as Cheng:2020ipp
$\mathcal{B}(D_{0}\to DP)$ $=$ $\frac{\tau_{D_{0}}p_{c}(m_{D_{0}},m_{D},m_{P})}{8\pi m^{2}_{D_{0}}}|A(D_{0}\to DP)|^{2},$ (46)
$\mathcal{B}(D^{*}\to DP)$ $=$ $\frac{\tau_{D^{*}}p_{c}^{3}(m_{D^{*}},m_{D},m_{P})}{6\pi m^{2}_{D^{*}}}|A(D^{*}\to DP)|^{2},$ (47)
$\mathcal{B}(D_{2}^{*}\to DP)$ $=$ $\frac{\tau_{D_{2}^{*}}p_{c}^{5}(m_{D_{2}^{*}},m_{D},m_{P})}{60\pi m^{2}_{D_{2}^{*}}}|A(D_{2}^{*}\to DP)|^{2},$ (48)
where the center of mass momentum $p_{c}(m_{D_{J}},m_{D},m_{P})\equiv\frac{\sqrt{\lambda(m_{D_{J}}^{2},m_{D}^{2},m_{P}^{2})}}{2m_{D_{J}}}$.
Four decay modes of the $D^{*}\to D\pi$ decays have been measured, and the data within $2\sigma$ errors are ParticleDataGroup:2022pth
$\mathcal{B}(D^{*0}\to D^{0}\pi^{0})$ $=$ $(64.7\pm 1.8)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(D^{*+}\to D^{0}\pi^{+})=(67.7\pm 1.0)\%,$
$\mathcal{B}(D^{*+}\to D^{+}\pi^{0})$ $=$ $(30.7\pm 1.0)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(D^{*+}_{s}\to D^{+}_{s}\pi^{0})=(5.8\pm 1.4)\%.$ (49)
Isospin violating decays $D^{*+}_{s}\to D^{+}_{s}\pi^{0}$ and later $D_{s0}^{0}\to D^{0}_{s}\pi^{0}$, which might decay via $\eta-\pi^{0}$ mixing Cho:1994zu , are not considered in this work. The experimental data of $\mathcal{B}(D^{*0}\to D^{0}\pi^{0})$ and $\mathcal{B}(D^{*+}\to D^{0}\pi^{+},D^{+}\pi^{0})$ will be used to obtain $\mathcal{B}(B\to D\pi\ell\nu_{\ell})$ with the $D^{*}$ resonances.
In the $D_{0}\to DP$ decays, only $D_{0}\to D\pi$ decays have been seen but have no data. Since $D_{0}\to D\eta$, $D_{0}\to D_{s}K$ and $D_{s0}\to DK$ are not allowed by the phase spaces, we assume $\mathcal{B}(D_{0}\to D\pi)=1$ to obtain four branching ratios of the $D_{0}^{0}\to D^{0}\pi^{0},D^{+}\pi^{-}$ and $D_{0}^{+}\to D^{+}\pi^{0},D^{0}\pi^{+}$ decays. And they are
$\mathcal{B}(D^{0}_{0}\to D^{0}\pi^{0})$ $=$ $(33.62\pm 0.03)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(D^{0}_{0}\to D^{+}\pi^{-})=(66.38\pm 0.03)\%,$
$\mathcal{B}(D^{+}_{0}\to D^{+}\pi^{0})$ $=$ $(33.18\pm 0.01)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(D^{+}_{0}\to D^{0}\pi^{+})=(66.82\pm 0.01)\%.$ (50)
The branching ratios of $D_{0}\to D\pi$ decays are the same as those in Eq. (50) when considering $D_{0}$ as the four quark state.
For the $D^{*}_{2}\to DP$ decays, since there is no experimental data of the branching ratios, we can not constrain $a^{D^{*}_{2}}_{01}$ directly. Nevertheless, $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ within $2\sigma$ errors are measured ParticleDataGroup:2022pth . The SU(3) relation of the decay amplitudes of the $D^{*}_{2}\to D^{*}P$ decays are given in Ref. Wan:2024ssl . Using $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ and assuming $\mathcal{B}(D^{*0}_{2}\to D\pi)+\mathcal{B}(D^{*0}_{2}\to D^{*}\pi)\leq 1$, $\mathcal{B}(D^{*+}_{2}\to D\pi)+\mathcal{B}(D^{*+}_{2}\to D^{*}\pi)\leq 1$ and $\mathcal{B}(D^{*+}_{s2}\to DK)+\mathcal{B}(D^{*+}_{s2}\to D^{*}K)\leq 1$, one can constrain the non-perturbative coefficients $a^{D^{*}_{2}}_{01}$, and we obtain that $|a^{D^{*}_{2}}_{01}|=25.14\pm 1.47$. Then the branching ratios of the $D^{*}_{2}\to DP$ decays can be predicted, which are given in the second column of Tab. 7. In addition, their decay width predictions and previous width predictions are also given in the third and fourth columns of Tab. 7, respectively. Our width predictions are about 1 time larger than ones in Ref. Wang:2014oca , nevertheless, they are very close to ones in Ref. Zhang:2016dom .
Table 6: The decay amplitudes for the $D^{*}_{2}\to DP$ decays under the SU(3) flavor symmetry.
     Decay modes Coupling vertexs      Decay modes Coupling vertexs
$D^{*0}_{2}\to D^{0}\pi^{0}$ $a^{D^{*}_{2}}_{01}/\sqrt{2}$ $D^{*+}_{2}\to D^{+}\pi^{0}$ $-a^{D^{*}_{2}}_{01}/\sqrt{2}$
$D^{*0}_{2}\to D^{0}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$ $D^{*+}_{2}\to D^{+}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
$D^{*0}_{2}\to D^{+}\pi^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}\to D^{0}\pi^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*0}_{2}\to D^{+}_{s}K^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}\to D^{+}_{s}\overline{K}^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{0}K^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{+}K^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{+}_{s}\eta$ $-a^{D^{*}_{2}}_{01}\Big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
Table 7: The branching ratio predictions of the $D^{*}_{2}\to DP$ decays within $2\sigma$ errors.
Decay modes Branching ratios (×$(\times 10^{-2})$ Decay widthes (MeV) Decay widthes from (MeV)
$D^{*0}_{2}\to D^{0}\pi^{0}$ $20.18\pm 1.77$ $9.59\pm 1.06$ $4.14^{+1.82}_{-1.57}$ Wang:2014oca , $12.0$ Zhang:2016dom
$D^{*0}_{2}\to D^{0}\eta$ $0.13\pm 0.03$ $0.06\pm 0.02$cdots
$D^{*0}_{2}\to D^{+}\pi^{-}$ $38.51\pm 3.41$ $18.30\pm 2.03$ $7.91^{+3.49}_{-3.00}$ Wang:2014oca , $22.8$ Zhang:2016dom
$D^{*0}_{2}\to D^{+}_{s}K^{-}$ $[1.82\times 10^{-15},6.56\times 10^{-6}]$ $[8.77\times 10^{-14},3.00\times 10^{-6}]$cdots
$D^{*+}_{2}\to D^{+}\pi^{0}$ $19.40\pm 1.63$ $9.33\pm 0.92$cdots
$D^{*+}_{2}\to D^{+}\eta$ $0.11\pm 0.04$ $(5.45\pm 2.15)\times 10^{-2}$cdots
$D^{*+}_{2}\to D^{0}\pi^{+}$ $39.88\pm 3.33$ $19.18\pm 1.90$cdots
$D^{*+}_{2}\to D^{+}_{s}\overline{K}^{0}$ $[1.26\times 10^{-14},1.80\times 10^{-6}]$ $[6.01\times 10^{-13},8.28\times 10^{-5}]$cdots
$D^{*+}_{s2}\to D^{0}K^{+}$ $42.32\pm 5.20$ $7.52\pm 0.95$ $3.35^{+1.48}_{-1.27}$ Wang:2014oca , $9.49$ Zhang:2016dom
$D^{*+}_{s2}\to D^{+}K^{0}$ $38.12\pm 4.72$ $6.79\pm 0.86$ $3.04^{+1.34}_{-1.15}$ Wang:2014oca , $8.61$ Zhang:2016dom
$D^{*+}_{s2}\to D^{+}_{s}\eta$ $0.58\pm 0.19$ $(10.23\pm 3.43)\times 10^{-2}$cdots
III.3 Numerical results of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
In terms of $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ given in Tabs. 5-5 and $\mathcal{B}(D_{J}\rightarrow DP)$ given in Eqs. (49-50) and Tab. 7, after considering the further experimental bounds of the resonant $B\to D\pi\ell^{+}\nu_{\ell}$ decays given in Eqs. (2-3) and (5-6), one can obtain the branching ratio predictions of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, and they are listed in the third, fourth and fifth columns of Tab. 2 for the $D^{*}$, $D_{0}$ and $D_{(s)2}^{*}$ resonances, respectively. Corresponding experimental data with $2\sigma$ errors are also listed in Tab. 2 for the convenience of comparison. Note that, since the vector resonances are also considered in this work, $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ in Eq. (1) and Eq. (4), which only conclude $D_{0}$ and $D^{*}_{2}$ resonances, are not used for our results. Many resonant branching ratios in Tab. 2 are predicted for the first time.
One can see that the the vector meson $D^{*}$ resonances give the dominant contributions in the $B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{0}\to\overline{D}^{0}\pi^{-}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$ decays, largely because of its proximity to the $D\pi$ threshold. Please note that decay amplitude of the $B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ decays is larger than ones of the $B^{+}\to D^{0}\pi^{0}\ell^{+}\nu_{\ell}$ decays by factor $\sqrt{2}$, nevertheless, the latter branching ratios are much larger than the former ones, since the most dominant resonance $D^{*0}$ cannot decay into $D^{-}\pi^{+}$ on its mass-shell ParticleDataGroup:2022pth . In previous studies, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*0}}=34.9\times 10^{-3}$ Cheng:1993ah , $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*-}}=16.7\times 10^{-3}$ Cheng:1993ah , $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*-}}=14.0\times 10^{-3}$ Kim:2017dfr , $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\tau^{+}\nu_{\tau})_{D^{*-}}=3.53\times 10^{-3}$ Kim:2017dfr , after considering the error, our corresponding results are consistent with them.
As for the scalar meson $D_{0}$ resonances and the tensor meson $D^{*}_{2}$ resonances, the experimental upper limit of $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{0}_{0}}$ gives further constraint on the $\mathcal{B}(B\to D\pi\ell^{\prime+}\nu_{\ell^{\prime}})_{D_{0}}$ predictions, and the experimental lower limit of $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*0}_{2}}$ gives further constraint on the $\mathcal{B}(B\to D\pi\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{*}_{2}}$ predictions. Our predictions for $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{-}_{0},D^{*-}_{2}}$ are more precise than their experimental measurements. The contributions of the $D_{0}$ and $D^{*}_{2}$ resonances are in the same order of magnitude in the $B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ and $B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. But the contributions of the $D_{0}$ resonances are larger than ones of the $D^{*}_{2}$ resonances in the $B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{+}\to D^{-}\pi^{+}\tau^{+}\nu_{\tau}$ and $B^{0}\to\overline{D}^{0}\pi^{-}\tau^{+}\nu_{\tau}$ decays.
The Belle II experiment has reported the branching ratios of the $B\to D\eta\ell^{\prime+}\nu_{\ell^{\prime}}$ decays with quite large errors, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{T}=(4.0\pm 4.0)\times 10^{-3}$ and $\mathcal{B}(B^{+}\to D^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{T}=(4.0\pm 4.0)\times 10^{-3}$ Belle-II:2022evt , which are not used for our predictions. From our predictions, one can see that the non-resonant branching ratios are dominant in the $B\to D\eta\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. And our predictions of $\mathcal{B}(B\to D\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ lie in the range of experimental data with $1\sigma$ error.
In addition, the interference terms between the non-resonant, the vector resonant, the scalar resonant and the tensor resonant contributions exist, and they might not be ignored if more than one kind of contributions are important in the decays, and they will be studied in our succeeding work.
All present experimental data of $\mathcal{B}(B\to DP\ell^{+}\nu_{\ell})$ may be explained by the SU(3) flavor symmetry approach. The SU(3) flavor breaking effects in the $B\to DP\ell^{+}\nu_{\ell}$ decays are in the similar to ones in the $B\to D^{*}P\ell^{+}\nu_{\ell}$ decays. As given in Tab. 2, the dominate SU(3) flavor breaking effects might appear in the non-resonant and the charmed tensor resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Nevertheless, there is only the data for $\mathcal{B}(B_{u,d}\to D\pi\ell^{+}\nu_{\ell})_{D^{*}_{2}}$, and there is not any data for $\mathcal{B}(B_{s}\to DK\ell^{+}\nu_{\ell})_{D^{*}_{s2}}$. Or there is only the data of $\mathcal{B}(B^{+}\to D_{s}^{-}K^{+}\ell^{+}\nu_{\ell})_{N}$, and there is no data of $\mathcal{B}(B_{u,d,s}\to DK\ell^{+}\nu_{\ell})_{N}$. Therefore, we can not directly judge how large the possible SU(3) breaking effects are in the $B\to DP\ell^{+}\nu_{\ell}$ decays.
Although the widths of all resonances are narrow, following Refs. Cheng:1993ah ; LeYaouanc:2018zyo , the width effects of $D^{*}$ mesons are analyzed. After considering the width effects of $D^{*}$ mesons, the decay branching ratios of $B\to D\pi\ell^{+}\nu_{\ell}$ are Cheng:1993ah ; Tsai:2021ota
[…]
where $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$ are obtained from to Eq. (28) and Eq. (47) by replacing $m_{D^{*}}\to\sqrt{t_{V}}$, respectively. There are two non-perturbative coefficients $V^{BD^{*}}(0)$ in $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $a_{01}^{D^{*}}$ in $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$. $V^{BD^{*}}(0)=0.65\pm 0.05$ from the data of $\mathcal{B}(B\to D^{*}\ell^{+}\nu_{\ell})$ listed in Tab. 5, and $a_{01}^{D^{*}}=8.42\pm 0.38$ from the data of $\mathcal{B}(D^{*+}\to D\pi)$ given in Eq. (49). $\Gamma_{D^{*0}}=(56.00\pm 5.74)\leavevmode\nobreak\ KeV$, which is obtained via the experimental data of $\mathcal{B}(D^{*0}\to D^{0}\pi^{0})$ in Eq. (49) and the SU(3) flavor symmetry in $D^{*}\to D\pi$. Following Refs. Cheng:1993ah , choosing $n=3$, the results of $\mathcal{B}(B\to D\pi\ell^{+}\nu_{\ell})_{D^{*}}$ are obtained, and they are listed in Tab. 2 by denoting $\dagger$. One can see that the results obtained by considering the $D^{*}$ width effects are slightly smaller than ones obtained by the narrow width approximation.
IV Summary
The semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonances, the vector resonances, the scalar resonances and the tensor resonances have been explored in terms of the SU(3) flavor symmetry based on the relevant experimental data. The amplitude relations of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the semileptonic $B\to D/D^{*}/D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays and the non-leptonic $D^{*}/D_{0}/D^{*}_{2}\to DP$ decays have been obtained, and then the resonant branching ratios have been obtained by the narrow width approximation after considering the resonant experimental data of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. Our main results can be summarized as follows.
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the central values of $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Other non-resonant decays are strongly suppressed by the narrow phase spaces.
For the charmed vector resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, they give the dominant contributions in the $B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{0}\to\overline{D}^{0}\pi^{-}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$ decays. Nevertheless, since the resonance $D^{*0}$ can not decay into $D^{-}\pi^{+}$, the total branching ratios of the $B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ decays are much smaller than ones of the $B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{0}\to\overline{D}^{0}\pi^{-}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$ decays.
As for the charmed scalar and tensor resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, our predictions of $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{D^{-}_{0},D^{*-}_{2}}$ are more precise than their experimental measurements. The contributions of the $D_{0}$ and $D^{*}_{2}$ resonances are in the same order of magnitude in the $B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ and $B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. But the contributions of the $D_{0}$ resonances are larger than ones of $D^{*}_{2}$ resonances in the $B^{+}\to\overline{D}^{0}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{0}\to D^{-}\pi^{0}\ell^{+}\nu_{\ell}$, $B^{+}\to D^{-}\pi^{+}\tau^{+}\nu_{\tau}$ and $B^{0}\to\overline{D}^{0}\pi^{-}\tau^{+}\nu_{\tau}$ decays.
Although the SU(3) flavor symmetry only can give the approximate predictions, they are still very useful for understanding these decays. Until now, our predictions of the $B\to DP\ell^{+}\nu_{\ell}$ decays are quite coincident with present experimental data, and our predictions could be tested in future experiments, such as LHCb and Belle II.
ACKNOWLEDGEMENTS
The work was supported by the National Natural Science Foundation of China (No. 12175088 and No. 12365014).